OpenStudy (anonymous):

can some one check my answer please! Determine whether each situation involves a permutation or a combination.Then find the number of possibilities. selecting nine books to check out of the library from a reading list of twelve 9! permutation

OpenStudy (dumbcow):

combination, you are picking a group of 9 books out of 12, order does not matter.

OpenStudy (dumbcow):

\[\frac{12!}{3!9!} = \frac{12*11*10}{3*2*1} = 2*11*10 = 220\]